OpenStudy (anonymous):

what is the greatest common factor of 28 and 42

OpenStudy (zzr0ck3r):

note we kept doing the same thing until we have no remainder, then we take the last remainder

OpenStudy (zzr0ck3r):

42=28*1+14 28=14*2 + 0 so 14

OpenStudy (calculusxy):